OpenStudy (anonymous):

What is the probability of choosing a 2 or a heart from a deck of cards? .00 .077 .25 .3077 What is the probability of choosing a 2 and a face card from a deck of cards .077 .25 .3077 .3269

OpenStudy (anonymous):

does this deck have jokers?

OpenStudy (anonymous):

or is it just standard 52

OpenStudy (anonymous):

always assume standard 52 on card problems... How many of the 52 cards are 2's or hearts ? 13 are hearts 4 are 2's 13 + 4 - 1 .....we minus 1 since we don't want to double count the 2 of hearts 16/52.....simplify.......4/13

OpenStudy (anonymous):

@guitar1329 standard 52

OpenStudy (anonymous):

@BangkokGarrett will the second one be 4/13 too 12 face cards + 4 2's= 16/52= 4/13